### v1

# Zero-shot forecasting of chaotic systems

Abstract
Time-series forecasting is a challenging task that traditionally requires specialized models custom-trained for the specific task at hand. Recently, inspired by the success of large language models, foundation models pre-trained on vast amounts of time-series data from diverse domains have emerged as a promising candidate for general-purpose time-series forecasting. The defining characteristic of these foundation models is their ability to perform zero-shot learning, that is, forecasting a new system from limited context data without explicit re-training or fine-tuning. Here, we evaluate whether the zero-shot learning paradigm extends to the challenging task of forecasting chaotic systems. Across $135$ distinct chaotic dynamical systems and $10^{8}$ timepoints, we find that foundation models produce competitive forecasts compared to custom-trained models (including NBEATS, TiDE, etc.), particularly when training data is limited. Interestingly, even after point forecasts fail, foundation models preserve the geometric and statistical properties of the chaotic attractors, demonstrating a surprisingly strong ability to capture the long-term behavior of chaotic dynamical systems. Our results highlight the promises and pitfalls of foundation models in making zero-shot forecasts of chaotic systems.
1 Introduction
Classical paradigms in machine learning (ML) require the model to be trained on data specific to the intended task. For example, to forecast the weather in Singapore, a model would need to be trained on past weather data from Singapore. However, recent work in statistical learning has highlighted the power of generative pre-trained models, which use probabilistic approaches and vast amounts of training data to build foundation models that can excel at diverse tasks without the need for separate retraining. In time-series forecasting, this paradigm shift has ignited an intense race to build general-purpose pre-trained models that can make zero-shot forecasts for any time series (Oreshkin et al., 2021; Garza & Mergenthaler-Canseco, 2023; Rasul et al., 2023; Jin et al., 2023; Gruver et al., 2024; Dooley et al., 2024; Liu et al., 2024; Woo et al., 2024; Ansari et al., 2024; Goswami et al., 2024). Such models have seen some initial success in forecasting real-world time series (Liang et al., 2024) but they have not been systematically tested on chaotic dynamical systems, especially in terms of their performance in long-term forecasting over an extended time horizon.
Figure 1: Chaos as a benchmark for zero-shot forecasting of time series. We use the $135$ chaotic systems from the dysts dataset (Gilpin, 2021) to generate chaotic trajectories from different initial conditions. Each trajectory is used to train the baseline deep-learning models (NBEATS, TiDE, etc.) and also provided as context to the pre-trained LLM (we use Chronos, a best-in-class foundation model for time series capable of zero-shot forecasting). Both the trained baseline models and Chronos are then asked to extend the trajectory by predicting the flow of the chaotic system. We measure the quality of the predictions in terms of both short-term accuracy and long-term attractor reconstruction. Across thousands of chaotic trajectories and over $10^{8}$ data points, we found that zero-shot forecasts from Chronos can be competitive in both short-term predictions and in capturing the long-term “climate” of the dynamics.
There are several reasons why such tests are interesting. First, to train foundation models for time series, the amount of high-quality time-series data needed is the single most crucial bottleneck. For this reason, a significant percentage of openly-available time-series data has been used to train these models. It is thus difficult to verify that the test set is not contaminated by time series related to those in the training set. In contrast, as far as we know, no trajectories generated by classical chaotic systems (e.g., Lorenz equations) have been used to train foundation models. Thus, time series from chaotic systems constitute an independent test set that can be used to quantify the generalization ability of foundation models. Second, chaotic dynamical systems have well-defined attractors that exhibit invariant statistical and geometric properties (fractal dimensions, Lyapunov exponents, power spectra, etc.). This allows us to quantify ML models’ ability to capture the long-term behavior of the system even after point forecasts inevitably fail (Pathak et al., 2018; Hess et al., 2023). Such tests are usually not possible for general time series. Third, the past few years have seen growing activities at the interface of physics and ML (Yu & Wang, 2024; Levine & Tu, 2024; Gilpin, 2024), with the cross-fertilization between ML and dynamical systems yielding advances in both directions (Weinan, 2017; Chen et al., 2018; Pathak et al., 2018; Li et al., 2020; Chen & Tao, 2021; Jordan et al., 2021; Gauthier et al., 2021; Levine & Stuart, 2022; Mikhaeil et al., 2022; Krishnapriyan et al., 2023; Yang et al., 2024). Benchmarking foundation models on chaotic systems introduces the possibility of applying dynamical systems techniques (e.g., Takens embedding theorem (Huke, 2006)) to understand the inner workings of these models and the origin of their generalization abilities.
In this paper, we set out to perform the first systematic evaluation of the zero-shot learning paradigm in the context of forecasting chaotic systems. A schematic summarizing our benchmark pipeline is presented in Fig. 1. We also show another schematic illustrating the difference between classical deep learning models and foundation models when making time series predictions (see Fig. 6 in the appendix).
Our study is also of intrinsic interest to scientific machine learning (SciML) and nonlinear dynamics communities. So far, the data-driven modeling approaches developed in these communities (reservoir computing (Pathak et al., 2018), PINN (Karniadakis et al., 2021), SINDy (Brunton et al., 2016), Koopman operators (Brunton et al., 2022), neural operators (Azizzadenesheli et al., 2024), etc.) still very much follow the classical train/test dichotomy. That is, to forecast the dynamics of the Lorenz oscillator, an ML model needs to be trained on data generated by the Lorenz equations. The model learns the chaotic dynamics by extracting the underlying vector field (or flow map) from time-series data during training. At first glance, it seems ludicrous that a model can effectively forecast chaotic dynamical systems without first explicitly learning the flow. A convincing demonstration of the possibility of zero-shot learning in a SciML context could lead to new forecasting tools and generate novel insights into chaotic systems.
On the more theoretical side of the inquiry, an emerging direction in SciML is to understand the out-of-distribution generalization ability of different data-driven modeling frameworks (Wang et al., 2020; Kong et al., 2021, 2023; Göring et al., 2024). This parallels a long line of research that investigates the generalization ability of neural networks (Neyshabur et al., 2018; Belkin et al., 2019; Baldassi et al., 2020; Xu et al., 2020; Feng & Tu, 2021; Nakkiran et al., 2021; Power et al., 2022; Liu et al., 2022c). For example, if a model was only trained on trajectories from a limited number of initial conditions, can it effectively extrapolate the learned dynamics to a different part of the phase space and forecast from a previously unseen initial condition (that is, far from any of the training initial conditions) (Zhang & Cornelius, 2023)? Foundation models that not only generalize to new initial conditions but also to new systems could introduce novel ideas and insights into this endeavor.
Our main contributions are:
  1. 1.
    A large-scale evaluation of the ability of time series foundation models to model physical systems outside of their training domain.
  2. 2.
    Discovery that foundation models produce zero-shot forecasts competitive with models custom-trained to forecast chaotic attractors. Moreover, larger foundation models produce better forecasts.
  3. 3.
    Observation of scaling of a foundation model’s zero-shot prediction ability with context lengths far exceeding typical correlation timescales of chaos, indicating in-context learning of chaotic dynamics.
  4. 4.
    Observation that foundation models retain long-term statistical properties of chaotic attractors, even after pointwise predictions fail.
2 Related work
Several works train transformers to perform long-horizon forecasting tasks (Li et al., 2019; Zhou et al., 2021, 2022; Liu et al., 2022b; Wen et al., 2022), obtaining leading results in long-horizon forecasting. However, recent works question their consistency and utility compared to properly-tuned simpler models (Lara-Benítez et al., 2021; Zeng et al., 2023; Das et al., 2023; Tan et al., 2024). Despite these debates, a unique property of large models like transformers is zero-shot generalization, in which they learn to perform a novel task without training the model weights on task-specific data (Brown, 2020). The resulting in-context learning strongly differs from prior approaches to forecasting chaotic systems, which focus on training the weights of models based on the past observed history of a system (Pathak et al., 2018; Gauthier et al., 2021; Vlachas et al., 2020). In-context learning has motivated the development of foundation models: large models pre-trained on vast amounts of data, which perform few-shot inference via prompting (Bommasani et al., 2021).
Several recent zero-shot forecasting models are modifications of large language models, which encode time series as tokens (Xue & Salim, 2023; Ansari et al., 2024; Gruver et al., 2024; Miller et al., 2024; Liu et al., 2024). Several of these models have been shown to exhibit in-context learning at test time (Lu et al., 2024; Gao et al., 2024; Liang et al., 2024).
Foundation models have recently been introduced for other scientific machine-learning tasks (Miller et al., 2024). These include models for partial differential equations (Yang et al., 2023; Rahman et al., 2024; Subramanian et al., 2024; Herde et al., 2024; Takamoto et al., 2022), numerical integration (Song et al., 2024), fluid flow prediction (Herde et al., 2024), molecular dynamics (Allen et al., 2024), weather forecasting (Nguyen et al., 2023; Bodnar et al., 2024), material discovery (Takeda et al., 2023), astrophysics (Parker et al., 2024), and electrocardiogram (ECG) analysis (McKeen et al., 2024). However, to the best of our knowledge, this work is the first large-scale evaluation of the zero-shot learning ability of foundation models on over $100$ chaotic systems, both in terms of short-term forecast accuracy and long-term attractor reconstruction performance.
3 A motivating example
Here, we chose Chronos (Ansari et al., 2024) to represent pre-trained models because it has been shown to outperform earlier foundation models for time series, such as TimeGPT and Moirai (Garza & Mergenthaler-Canseco, 2023; Woo et al., 2024). Chronos internally uses a large language model based on the text-to-text T5 transformer model family (Raffel et al., 2020). It introduces a scaling and quantization layer, which converts continuous-valued univariate time series into a set of discrete tokens, with vocabulary size acting as a model hyperparameter. The model was trained on diverse time series spanning $\sim\!10^{11}$ observations drawn from $42$ synthetic and real-world settings, but the training data does not contain any dynamical systems. We evaluate five pre-trained variants of Chronos, denoted by the sizes of the underlying T5 architecture: $8M$, $20M$, $46M$, $200M$, and $710M$ parameters.
Figure 2: Chronos can make meaningful zero-shot forecasts of chaotic systems. Here, we use Chronos to predict the $x$ and $y$ components of the Lorenz oscillator. Since Chronos makes probabilistic forecasts, we run the model independently $20$ times and plot the mean forecast. We note that there is very little variation between different runs, implying that the uncertainty of the forecast is low. The zero-shot forecasts match remarkably well with the ground truth for both short-term prediction and long-term attractor reconstruction.
Fig. 2 shows zero-shot forecasting of the Lorenz oscillator (defined in the appendix), a well-studied chaotic dynamical system, using Chronos-200M. The only data Chronos has access to are the $512$ data points that serve as the context for the prediction (pink curve from $t=1$ to $t=512$). Because Chronos is a univariate forecast model, we separately forecast each coordinate of the attractor and reset the model state between each forecast. Forecasting chaotic systems based on partial observations (e.g., only having access to the $x$ or $y$ coordinate of the Lorenz oscillator) is an extremely difficult task in nonlinear dynamics (Ratas & Pyragas, 2024). Despite this challenge, the prediction closely tracks the ground truth for over $100$ time steps and, even after diverging from it due to chaos, remains in the vicinity of the strange attractor.
Interestingly, although Chronos predicts $x$ and $y$ separately, it maintains a positive correlation between $x$ and $y$ so they have the same sign most of the time (which is necessary for accurately reconstructing the attractor). This suggests that Chronos internally models $y$ when forecasting $x$ and vice-versa. In nonlinear dynamics, this process is possible due to Takens’ theorem, which states that low-dimensional measurements can reveal unmeasured dynamical variables using delay embedding (Huke, 2006).
However, the performance of Chronos, while impressive, can also be fragile. Keeping everything unchanged, simply starting the context trajectory from a different initial condition on the attractor can significantly degrade the accuracy of Chronos’s prediction (see Fig. 9 in the appendix). So how good is Chronos at forecasting chaotic systems, truly? More generally, is zero-shot forecasting from foundation models a promising alternative to custom-trained models when it comes to predicting chaotic systems? To answer these questions, we next perform systematic benchmarks that average over a diverse set of chaotic systems and different initial conditions.
4 Methods
4.1 Chaos as a forecasting benchmark
A natural test of modern forecasting methods comes from the field of nonlinear dynamics, which seeks to quantify the degree to which systems are predictable. A key measure is the largest Lyapunov exponent $\lambda$, an invariant property associated with every set of differential equations that quantifies the rate at which small errors accumulate. Systems that approach a periodic orbit or an equilibrium exhibit zero or negative Lyapunov exponents because different initial conditions converge to the same state. In contrast, chaotic systems exhibit positive Lyapunov exponents, implying that small changes in the initial conditions or the model parameters lead to trajectories that (at least initially) diverge exponentially. When modeling such systems, a small error will compound over a characteristic timescale, the Lyapunov time, $\tau\equiv\lambda^{-1}$, making highly-chaotic systems (those with small $\tau$) difficult to forecast.
The dysts dataset represents a standardized benchmark of $135$ low-dimensional chaotic systems, described by ordinary differential equations that have been aligned with respect to their dominant timescales and integration steps (Gilpin, 2021, 2023). Each equation in the database has its own Lyapunov time $\tau$, as well as distinctive chaotic dynamics annotated by mathematical properties such as entropy or fractal dimension. The dataset can be re-integrated to produce chaotic multidimensional time series of arbitrary length, granularity, and noise level. Here, in order to match the typical properties of the real-world time series used to train Chronos, we generate a new dataset of similar time series. For each chaotic system, we integrate a set of trajectories using an implicit Runge-Kutta scheme, and then downsample them to a uniformly coarse granularity of $30$ timepoints per Lyapunov time $\tau$. However, we observe only a weak dependence of our forecast results on the data granularity (Appendix).
4.2 Baseline experiments
Our baseline experiment design matches prior works (Gilpin, 2021, 2023; Godahewa et al., ; Schötz et al., 2024). For each of the $135$ chaotic dynamical systems, $20$ trajectories of length $812$ are generated, each originating from a random initial condition on the attractor. This produces a set of $2700$ ($135\times 20$) multivariate time series, which have dimensionalities between $3$ and $6$ depending on the particular dynamical system. All time series are then split into training sets consisting of the first $512$ points of each time series, with the last $300$ timepoints set aside to determine final test scores. For experiments with varying context lengths, trajectories are extended backwards in time, so that the $300$ test points remain the same.
For the baseline models, hyperparameter tuning is performed separately for each of the $135$ dynamical systems. For a given dynamical system, each of the $20$ training trajectories is divided into a true training set comprising the first $435$ timepoints, and a validation set of the last $77$ timepoints. For each set of hyperparameters, a model is trained on the true training set and then evaluated on the validation set. The validation scores are averaged over the $20$ trajectories, and the hyperparameters from the best-performing model are selected. A model is then initialized with those hyperparameters, and it is trained on the full $512$ timepoints. The model is then tasked with autonomously generating a forecast of the next $300$ timepoints (around $10$ Lyapunov times), which are compared against the ground-truth trajectories to generate overall model scores. The testing dataset is therefore causally disconnected from the training data at all times.
To match the design of Chronos, for multivariate dynamical systems, each baseline model is separately trained and tested along each dimension, and the results are averaged. This channel-independent forecasting task is intrinsically harder than providing full state information, because the models cannot leverage the mutual information between different dimensions (Ratas & Pyragas, 2024). However, recent works on large-scale forecast models actually obtain stronger results by isolating input channels, because the resulting model class is more expressive (Nie et al., 2023). We thus do not expect Chronos’s performance to improve if it were instead trained to produce multivariate forecasts (i.e., one in which $x$,$y$,$z$ are jointly embedded and tokenized).
The experiments yield $2700$ distinct forecasts of $300$ timepoints each along $3-6$ dimensions depending on the underlying chaotic system, all generated by separately-trained forecast models. Our large-scale experiments thus span $5.5\times 10^{7}$ training points, $3.2\times 10^{7}$ test points, and $3.2\times 10^{8}$ generated forecasts across all models. The experiments require $10^{4}$ walltime compute hours on an Nvidia A100 GPU.
Our baseline models include NBEATS (Oreshkin et al., 2019), a hierarchical neural network model that has been shown to perform particularly well on dynamical systems forecasting tasks (Gilpin, 2021, 2023). TiDE (Das et al., 2023), a recent model that addresses several known computational limitations of Transformer class models on forecasting time series. A next-generation reservoir computer (NVAR) (Gauthier et al., 2021), which has a strong inductive bias for learning dynamical systems and which has previously been found to perform well on chaotic systems (Gilpin, 2023). We also include a small encoder-decoder Transformer with $0.5M$ trainable parameters, as well as an LSTM (Vaswani et al., 2017; Hochreiter, 1997).
In principle, the baseline models have a wide variety of additional hyperparameters available to tune, such as optimizer settings, reservoir or recurrent layer size, etc. Here, we focus on the lookback window, which is a common hyperparameter across all forecast models. It is also analogous to the context window in Chronos, for which we tune no other hyperparameters in the zero-shot setting.
4.3 Metrics
We use four metrics to evaluate forecasting quality: Symmetric Mean Absolute Percentage Error (sMAPE).
$\text{sMAPE}(\mathbf{x},\hat{\mathbf{x}})\equiv 2\dfrac{100}{T}\sum_{t=1}^{T}\dfrac{\left|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\right|}{\left|\mathbf{x}_{t}\right|+\left|\hat{\mathbf{x}}_{t}\right|},$
where $\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}$ correspond to the true test values of a time series up to a maximum forecast horizon $T$, and $\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2},...,\hat{\mathbf{x}}_{T}$ are the predictions of a forecast model at those same timepoints. Prior studies have evaluated the suitability of various error metrics in evaluating forecast accuracy (Hyndman & Koehler, 2006; Makridakis et al., 2022), including specifically on dynamical systems prediction (Gilpin, 2021, 2023), and found that sMAPE strongly correlates with other metrics (e.g. RMSE, NRMSE, MASE, Spearman correlation) while exhibiting favorable properties like a bounded range.
Valid Prediction Time (VPT). The first forecast horizon at which the sMAPE exceeds a fixed threshold $\epsilon$ (Vlachas et al., 2020).
$\text{VPT}\equiv\text{argmax}_{t_{f}}\{t_{f}|\text{sMAPE}(\mathbf{x}_{t},\hat{\mathbf{x}}_{t})<\epsilon,\;\forall t<t_{f}\}.$
We set $\epsilon=30$, as in prior studies (Vlachas et al., 2020; Gilpin, 2023).
Correlation Dimension ($d_{\text{frac}}$). For chaotic dynamical systems, the long-term distribution of observed data points approximates a fractal object known as the strange attractor. Fractals have space-filling properties that are intermediate between integer dimensionalities, and every strange attractor has a unique and invariant fractal dimension. The correlation dimension non-parametrically estimates the fractal dimension from a time series, by calculating the scaling of the number of other attractor points that fall within a given radius of each point (Grassberger & Procaccia, 1983). We compute the correlation dimension using all data points from a model’s forecasts and report the root mean square error between the inferred correlation dimension and the ground truth.
Kullback–Leibler Divergence between attractors ($D_{stsp}$). We compute the KL Divergence between the original and reconstructed attractors, following previous works (Hess et al., 2023; Göring et al., 2024). To perform the computation, we center a Gaussian distribution at each point from the true and reconstructed trajectories. We then use a sampling-based approach to estimate the KL Divergence between these Gaussian mixtures (Hershey & Olsen, 2007). This metric measures whether two attractors have matching distributions, and it largely agrees with the correlation dimension. We thus report the KL Divergence results in the Appendix.
5 Results
5.1 Comparing zero-shot models and fully-trained models.
To evaluate the effectiveness of zero-shot forecasting for chaotic systems, we evaluate the performance of Chronos and the baseline models on the dysts benchmark (Fig. 3). Across the $135$ systems, the median VPT of the three largest zero-shot Chronos models is statistically indistinguishable, while the smaller models exhibit significantly smaller VPT ($p<10^{-3}$, non-parametric Friedman test, $N=135$). Scaling of performance with model size indicates that the larger models exhibit better generalization properties, because the chaotic systems dataset strongly differs from their training data. This finding supports the premise of the foundation model paradigm for chaotic systems, because it shows that the sheer scale of a domain-agnostic model, when matched with sufficient training, improves forecasts. Compared to the fully-trained baseline models, the three largest zero-shot forecast models outperform all except for NBEATS (Friedman, $p<10^{-3}$, $N=135$). While recurrent neural networks and next generation reservoir computers have previously shown promising forecast results for dynamical systems (Vlachas et al., 2020; Gilpin, 2021; Gauthier et al., 2021), they underperform zero-shot models in the data-limited setting investigated here. However, when given enough training data, it has been shown that these models can achieve longer prediction horizons (Gauthier et al., 2021; Gilpin, 2023; Pathak et al., 2018). In contrast, the context length of Chronos and other attention-based forecast models is limited, and they are most effective when data is scarce.
Figure 3: Zero-shot models of chaotic systems are competitive with custom-trained models. Zero-shot forecasts from Chronos for five different model sizes (left), compared to other forecast models directly trained on the points given to Chronos as context (right). Inset plots show the valid prediction times (VPT), the first time each forecast exceeds an error limit. All error bars are over $135$ chaotic systems, each with $20$ distinct initial conditions.
We emphasize that the similarity of the error curves in Fig. 3 does not arise from a lack of sensitivity in the forecast metrics. When the baseline models are instead given full state information (multivariate forecasting), the prediction task becomes easier, resulting in lower sMAPE and higher VPT across all systems (see Appendix). These results underscore that partial observability, which characterizes most practical forecasting tasks (Ratas & Pyragas, 2024), is quantifiably harder for current forecasting models. The zero-shot models perform nearly as well as state-of-the-art, fully-trained models in this setting, reaching a VPT as high as 1 Lyapunov time.
Historically a prediction time of 1 Lyapunov timescale has been considered prohibitive even for fully-trained forecast models. This is because both observational and modeling error compound over this timescale (Palmer, 2000; Medio & Lines, 2001). Chronos’s ability to consistently forecast up to 1 Lyapunov time, without prior training on dynamical systems, suggests the advantages of its large-scale training on diverse time series. This scale allows it to extract generic predictive features from time series, which also prove to effectively represent nonlinear dynamics. A similar concept occurs in computer vision, in which convolutional neural networks tend to learn generic Gabor-like feature extractors in early convolutional layers (Zeiler & Fergus, 2014). The ability of Chronos to generate meaningful forecasts suggests that these learned nonlinear features, coupled with high dimensionality both in the input feature space (context length) and internal model dynamics, mitigate the intrinsic chaoticity of the underlying systems. In dynamical systems theory, recent works on Koopman operators emphasize that appropriately-selected nonlinear transformations make chaotic systems appear more linear (and thus predictable) in higher-dimensional spaces (Mezić, 2013; Brunton et al., 2022). As Chronos contains tens of millions of internal weights, it has an intrinsic advantage due to its scale, which counteracts its low inductive bias when compared to forecasting models specialized in dynamical systems, such as next-generation reservoir computers.
Next, we quantify Chronos and the baseline models’ ability to capture the long-term behavior of chaotic systems after point forecasts inevitably fail. This corresponds to a global measure of forecast quality: how well does a model capture the shape of the strange attractor and reproduce the statistics of major dynamic events, even if not necessarily their particular timing? In forecasting, this problem is known as predicting the climate, rather than the weather (Patel et al., 2021; Bramburger & Fantuzzi, 2024).
Figure 4: Zero-shot forecast models effectively capture attractor geometry. (A) Example forecasts produced by the zero-shot and trained models, for $20$ initial conditions from the Lorenz chaotic attractor. (B) The correlation between the fractal dimension of the predicted attractor and the true attractor (Spearman’s rank-order coefficient, $N=2420$ points, $p<10^{-3}$ for all cases), versus the VPT of the corresponding model. The red markers represent variants of Chronos with different model sizes: tiny ($8M$ parameters), mini ($20M$), small ($46M$), base ($200M$), and large ($710M$). The blue markers represent the baseline models. Models closer to the top capture the attractor geometry better and models closer to the right make accurate point forecasts for longer. Error bars are standard errors over $135$ dynamical systems, each with $20$ different initial conditions.
Figure 4 plots the correlation dimension accuracy (which measures long-term attractor reconstruction quality) against the VPT (which measures short-term forecast quality) for different models. NBEATS performs the best in both metrics, likely because very high pointwise accuracy necessarily leads to high global attractor quality. Generally, this trend holds across both zero-shot models and baseline models. However, within each model class a few surprises emerge: the fully-trained small Transformer, which produced relatively weak forecasts, captures the attractor shape as accurately as the zero-shot models. This observation suggests that attention-based models, which process their entire context simultaneously, have an innate advantage in capturing the long-term structure of attractors—a finding that mirrors similar results for language models (Brown, 2020). Consistent with this interpretation, we observe weak attractor reconstruction accuracy from the LSTM and NVAR models, which both operate sequentially and downweight earlier parts of their context. To ensure that these results are not a consequence of our choice of metric, we also evaluated attractor reconstruction quality using the KL Divergence between the true and forecasted attractors, and we found the same trends (see Appendix).
5.2 Accuracy and cost of in-weights versus in-context learning for varying data sizes.
We next evaluate how variation in context length affects the performance of Chronos. We vary the context length of the base Chronos model between $5$ and its maximum value of $512$ and repeat our zero-shot forecast experiments. We also select the best-performing traditional model, NBEATS, and fully train it (including cross-validation to set the lookback window) over the same points given to Chronos as context. We find that the VPT of Chronos increases monotonically with context length, even as it reaches over $17$ Lyapunov times. This timescale extends well-beyond the $\sim\!\!1$ Lyapunov timescale over which chaotic systems typically become decorrelated ($\langle\mathbf{x}(t)\mathbf{x}(t+\tau)\rangle_{t}=0$) (Shaw, 1981). This regime also exceeds the typical range of Takens’ embedding theorem, because time series are usually lifted using delay embeddings over timescales $<\tau$.
Chronos’s increasing forecasting performance stems from its ability to perform in-context learning, in which additional context points from an attractor act analogously to prompting in language models (Brown, 2020; Li et al., 2023). This mechanism underlies our earlier observation that the model’s generalization ability improves with its size. While early points in a long context are decorrelated with the predictions, they are drawn from the same underlying distribution. Chaotic attractors are ergodic, and so they have a well-defined stationary distribution of expected states known as their natural measure (Ott, 2002). As in language prediction, the early context points provide information about the measure and the conditional probabilities of successive data points (Xie et al., 2022). A similar phenomenon is observed for reservoir computers, a type of recurrent neural network in which a trainable linear layer is applied to a random transformation of the context data (Jaeger & Haas, 2004; Pathak et al., 2018). In this setting, extended context allows the reservoir to gradually synchronize with the dynamical system being learned (Lu & Bassett, 2020).
We next consider the practical question of whether the foundation model paradigm—pretraining on domain-agnostic settings, and then specializing to a task—confers computational advantages over directly training a smaller model from scratch. We measure the walltime of training and inference on a single A100 GPU node  [^1] Walltime is an imperfect measure of computational costs, because different models are specialized for different hardware (such as GPU acceleration or parallelization). Nonetheless, walltime within a given model class provides a proxy for a model’s performance in real-world settings.. We find that the computational cost of Chronos can be favorable at long context lengths when compared to NBEATS (Fig. 5B).
Figure 5: Effect of context length on performance and costs. (A) The Valid Prediction Time (VPT) for base Chronos as the context length is increased, and the VPT of NBEATS tuned and trained on the same data. For these experiments, NBEATS was trained in a multivariate setting (channel-mixing), giving it a slight advantage over Chronos. (B) The single-node walltime for zero-shot forecasts with Chronos-base, compared to the training (including hyperparameter tuning) and inference costs of NBEATS. All error ranges denote standard deviations over $135$ dynamical systems, each with $20$ different initial conditions. .
The inference time of Chronos is bounded by the quadratic scaling of attention layers with the context length. This limitation motivated newer architectures like Mamba (for language) and TiDE (for time series), which exhibit linear scaling. However, despite the relatively slow inference at small context windows, we find that Chronos can be very efficient when working with long context, making it a viable choice for many practical applications. In terms of the prediction horizon, Chronos exhibits the same linear scaling of cost as auto-regressive models (RC, LSTM, NVAR, etc.).
6 Conclusion and future directions
We have performed the first large-scale evaluation of zero-shot forecasting models on the classical problem of predicting chaos. Our most striking finding is that a large pre-trained model can successfully forecast chaotic systems for up to one Lyapunov time, beyond the expected degree of predictability, even though it was not directly trained on chaotic systems. The resource requirements of inference-only zero-shot forecasting are negligible compared to fully training deep-learning models such as NBEATS, particularly when long context lengths and lookback windows are used. Moreover, zero-shot models perform well without hyperparameter tuning. All in all, the success of Chronos indicates that many aspects of chaotic dynamics can be captured by generic high-dimensional transformations, suggesting that the internal representations used by Chronos to learn dynamical systems may provide insight into other time series tasks, such as system identification or bifurcation detection. It also supports the hypothesis that there is a common “language” for time series—universal features and structures shared by time series across different domains that make transfer learning possible.
On a conceptual level, unpredictability in chaotic systems arises from the rapid growth of the gap between the true trajectory and its approximation by a forecast model—motivating the intuition that Lyapunov time bounds predictability. The long lookback of models like Chronos allows them to leverage information from multiple past timepoints, and thus stabilize accumulation of error relative to forecast models that only consider the most recent timepoints (Viswanath, 2001). In this sense, long-context forecasting resembles multistep integration (Zhang & Cornelius, 2023, 2024). Recent work on dynamic mode decomposition and Koopman operator inference take this idea even further, by showing that time delays can lift dynamical systems into spaces where the dynamics are nearly linear (Brunton et al., 2017). We therefore broadly interpret the zero-shot capabilities of Chronos, which improve with model size, as illustrating the intrinsic inductive bias that comes from lifting nonlinear time series to very high dimensions. However, this does not fully explain our observation that long context windows, spanning multiple Lyapunov times, improve zero-shot forecasts. Instead, we attribute this phenomenon to the recent discovery of in-context learning in pre-trained forecast models, which is only recently starting to be explored in SciML Yang et al. (2023); Subramanian et al. (2024).
Our study therefore evaluates the suitability of the foundation model paradigm for SciML tasks. An important future direction for our investigation is task-specific tuning, in which the weights of large pre-trained models like Chronos are fine-tuned on a small number of example chaotic time series. This differs from the zero-shot in-context learning that we discuss above, and recent foundation models for partial differential equations have found that in-weights tuning can improve generalization (Subramanian et al., 2024). In our initial experiments, we found that very long time series, at least two orders of magnitude greater than the context lengths used here, were required to stably update the weights of Chronos and change validation scores. This mirrors results for large language models, where in-context learning has been shown to be preferable when few examples of the target task are available (Liu et al., 2022a).
7 Acknowledgments
YZ acknowledges support from the Omidyar Fellowship and NSF DMS 2436231. WG was supported by NSF DMS 2436233. This project has been made possible in part by Grant No. DAF2023-329596 from the Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation. Computational resources for this study were provided by the Texas Advanced Computing Center (TACC) at The University of Texas at Austin.
8 Code Availability
All zero-shot benchmark forecast results and scripts are available online at https://github.com/williamgilpin/dysts. The dynamical systems benchmark dataset is available online at https://github.com/williamgilpin/dysts_data.
Appendix A Additional schematics
Figure 6: Difference between baseline models and foundation models in forecasting chaotic systems. Classical deep-learning models (i.e., baseline models) forecast a chaotic system by learning the underlying vector field or flow map. To achieve this, the model adjusts its weights based on data from the same chaotic system. In contrast, foundation models (e.g., Chronos) do not train directly on the system they want to predict. Instead, they aim to “learn the language of time series” (Ansari et al., 2024) by training on vast amounts of time series data from diverse domains. After that, foundation models can make zero-shot forecasts on any (previously unseen) chaotic system based on a short context trajectory.
Appendix B Additional Methods
B.1 Lorenz equations.
Lorenz oscillator is one of the most studied chaotic systems and is described by the following equations:
$\begin{split}\dot{x}&=\sigma(y-x),\\ \dot{y}&=x(\rho-z)-y,\\ \dot{z}&=xy-\beta z,\end{split}$
where the default parameter values are $\sigma=10$, $\rho=28$, and $\beta=8/3$.
B.2 Measuring attractor similarity.
We measure attractor similarity using an approach introduced in previous works (Hess et al., 2023; Brenner et al., 2022). The state space divergence between the true and generated attractors is given by the Kullback-Leibler divergence between the distributions $p(\mathbf{x})$ and $q(\mathbf{x})$,
[…]
$D_{stsp}\approx\frac{1}{n}\sum_{i=1}^{n}\log\left(\frac{\hat{p}(\mathbf{x}^{(i)})}{\hat{q}(\mathbf{x}^{(i)})}\right),$
where $\mathbf{x}^{(i)}$ are samples drawn from the true orbit (Hershey & Olsen, 2007).
Appendix C Additional experiments
Figure 7: Zero-shot forecast models capture attractor geometry well, as measured by the KL Divergence. The state space divergence $D_{stsp}$ between the predicted attractor and the true attractor, versus the VPT of the corresponding model. The red markers represent variants of Chronos with different model sizes: tiny ($8M$ parameters), mini ($20M$ parameters), small ($46M$ parameters), base ($200M$ parameters), and large ($710M$ parameters). The blue markers represent the baseline models. Models closer to the bottom capture the attractor geometry better, and models closer to the right make accurate point forecasts for longer.
Figure 8: Comparison of univariate versus multivariate baseline forecasts of chaotic systems. Because Chronos is a univariate forecast model that predicts each time series channel independently, the baseline experiments we present in the main text (left panel here) involve channel-independent training, in which each baseline model is separately trained and tested on each dimension of the input time series. We repeat these experiments in a multivariate setting, by retraining the baseline models simultaneously on all dimensions (right panel). All error bars are over $20$ distinct initial conditions for each of the $135$ chaotic systems.
Figure 9: Chronos performance can be sensitive to initial conditions. Here, we use Chronos to predict the $x$ and $y$ components of the Lorenz oscillator. Everything is kept the same as in Fig. 2 except the initial condition used to generate the context trajectory. Compared to Fig. 2, we can see the Chronos prediction becomes noticeably worse.
Figure 10: Correlation between forecasts and invariant properties. The correlation between the Lyapunov exponent of each of the $135$ chaotic systems, and the sMAPE error of a forecast model, as a function of the prediction horizon.
Figure 11: Zero-shot attractor reconstruction accuracy scales with model size. The Spearman correlation between the fractal dimension of Chronos’s predictions, and the true fractal dimension of the underlying system, compared to the number of trainable parameters in the Chronos model.
Figure 12: Naive forecasts underperform all models evaluated. The growth of the sMAPE error for a naive constant forecast, in which the most recent training point is carried forward as the prediction for all future values. The shaded region corresponds to standard error across $135$ dynamical systems, with $20$ initial conditions each..